OpenStudy (anonymous):

Given: PA is tangent to the circle with center O. Prove: triangle APB is similar to triangle CPA. A, B, and C are points on the circle. P is outside the circle.

OpenStudy (anonymous):



Directrix (directrix):

  Given: PA is tangent to the circle with center O. Prove: triangle APB is similar to triangle CPA. Statements 1. Segment PA is tangent to the circle with center O. 2. m<CAP = (1/2)* m Arc AC 3. m<B = (1/2) * m Arc AC 4. <CAP ≅ < B 5.< P ≅ <P 6. Triangle APB ~ Triangle CPA

Directrix (directrix):

@kleduff87 Your task is to complete the reasons. Post what you get. Okay?